A copper microstrip transmission line is to be designed for a 100 Ξ© characteristic impedance. The substrate is 0.51 mm thick, πœ€π‘Ÿ = 2.2 and tan Ξ΄ = 0.001.

At 5 GHz, find:

3.1.1. The guide wavelength on the line (in cm) and;

3.1.2. The total attenuation in dB/cm.


1
Expert's answer
2022-03-01T17:32:51-0500

"\lambda=\frac {c}{\nu\sqrt{e_r}}=8.4~cm,"

"A=\gamma R\frac L{\sin\phi}=64~\frac{dB}{cm}."